# Responsibility and Engagement – Evaluating Interactions in Social Robot Navigation

### v1

@@ I Introduction
One goal of Social Robot Navigation (SRN) research is to enable robots to seamlessly navigate complex, unstructured, and potentially crowded environments. Interactions between robots and humans are at the core of SRN. Studies show that humans interacting with publicly deployed robots often compensate for the robots’ lack of navigation capabilities, thereby taking a major share of preventing collisions and conflicting situation [1, 2]. While this might be acceptable for the occasional robot, it could turn out to be a nuisance once more robots are deployed in public. Therefore, research on socially acceptable planning algorithms as well as suitable evaluation metrics has become increasingly important [3, 4, 5]. To assess the quality of a navigation algorithm, a meaningful quantitative evaluation of the output trajectories is paramount. Such metrics can also be part of target functions for learning-based approaches. In this work, we focus on metrics which evaluate various aspects of social compliance of motion behavior, i.e., trajectories. Recent work [6] pointed out that current metrics fail to address the causal component in interactions between two agents. The degree to which an agent contributed to resolving the conflict is vital for evaluating the quality of said agent’s behavior. To this end, the authors proposed a ”Responsibility” metric aiming to quantify the share that agents contributed to de-escalating a potential conflict.
@@ II Related work
Recent work [6] showed that existing metrics do not capture important aspects of two moving agents interacting with each other. Specifically, the authors analyze kinematics-based metrics (e.g. an agent’s maximum acceleration or mean jerk), metrics based on the distance between two agents (e.g., Clearing Distance or Space Compliance Rate) and prediction-based metrics like Minimum Time-To-Collision or Projected Path Duration [7, 3] on a simple scenario: two agents of varying capabilities (oblivious or foresighted) are facing each other on a head-on collision course. While all of the established metrics show meaningful information, they fail to fully highlight a key information: which of the agents (if any) were causally responsible for avoiding the conflict. [6] propose the Responsibility metric, which is based on predicted minimal distances. In the presented scenarios, the Responsibility metric shows which agent’s behavior contributed to the relaxation of the conflict. However, the set of tested scenarios is limited, and all scenarios start with both agents being at full conflict already. Furthermore, no scenarios with more than two agents are addressed.
Other work has also addressed the topic of responsibility in the context of navigation explicitly. [8] captures causal responsibility of moving agents in their proposed FeAR metric. Similar to [6], the authors assume that attribution of responsibility should be related to actions (or their lack of). The proposed approach is, however, based on a discrete, grid-based representation, and evaluated on simplified road traffic scenarios with strong assumptions about lanes and intersections. Hence, FeAR is difficult to apply to continuous motion data in unstructured environments.
@@ III Responsibility and Engagement Metrics
This section extends the Responsibility metric presented in [6]. First, we revisit the calculation of the underlying conflict potential in Section III-A. We then propose to derive the conflict measure by considering the previously unaddressed conflict buildup phase using a normalization step in Section III-B. Section III-C explains how the agents’ actions are taken into consideration for the escalation or de-escalation of the conflict, before Section III-D shows the updated Responsibility metric and introduces the Engagement metric. Program code for calculating the Responsibility and Engagement metrics from trajectories can be found in a GitHub repository111Will be disclosed after double-blind review..
@@ III-A Conflict Potential
We use the definition of [6] to calculate the Conflict Potential between two agents (referred to as agent 1 and agent 2) in an interaction. Here, the Distance at Closest Encounter (DCE) [9] of two trajectories plays a central role. To calculate the Conflict Potential, a predictive DCE (pDCE) is used, which captures the predicted proximity of two agents continuing along their current trajectories with constant velocity. Figure 1 shows how the pDCE captures the minimum distance between two agents when their movements are predicted based on a constant velocity assumption.
@@
where and are the radii of agents 1 and 2, assuming circular agents.
In a nutshell, the conflict potential equals one if the agents are on a direct collision course. It decreases as their predicted center positions at the point of closest encounter move further apart, becoming zero when they are predicted to just miss each other narrowly.
@@ III-C Agent Behavior and Conflict Contribution
Typically, time diminishes an existing conflict if the agents are still in conflict, but distance between them is increasing. One interpretation is that is the residual conflict which was not fully resolved by the agents.
@@ IV-B Experiment 1: Dyadic Interactions
This experiment provides two insights: First, since the Responsibility metric is based on the pDCE, it covers diverse interactions, regardless of the orientation of both agents toward each other. Also, it captures both longitudinal and lateral behavior changes. Second, while Responsibility is a valuable metric to investigate behavior in interactions, it does not entail any direct normative interpretation. In order to evaluate if a behavior was acceptable, the context is vital. Consider Scenarios 2 and 3, where only one agent is responsive, for the Oncoming and Overtaking case. Normatively, we would expect an agent to be alert, taking others into account when planning its trajectory. Hence, a Responsibility of zero typically corresponds to suboptimal (or at least unfriendly) behavior. However, in Scenario 3 of the Overtaking geometry, Alice’s oblivious behavior is fully acceptable, since she can’t be expected to see the robot approaching from behind. Vice versa, in Scenario 2, the robot is threatening to run over Alice from behind - which is completely unacceptable, since Alice has to jump away to save herself. Similarly, in the Following case, Alice’s Engagement would rise if she slowed down while the robot is approaching from behind (since she actively changes her behavior).
@@ IV-E Experiment 4: Playing Catch
Figure 8 shows one interaction of the game. The robot starts on the bottom right corner going diagonally to the upper left corner, with Alice starting in the bottom left corner, tying to intercept its way. The robot reacts by evading to the left. Alice readjusts the course but overshoots, starting to go in a narrow right curve, just avoiding a collision. the robot keeps evading. Since this point in time marks the time of closest encounter (TCE, see Eq. 3), the rest of the interaction (i.e., Alice and the robot completing the circle) is not captured.
@@ IV-F Experiment 5 - Personal Space Compliance
In Eq. 2, we can choose the agent sizes to determine which predicted distance leads to a potential conflict. Accordingly, it is possible to evaluate and with respect to proxemics [13] instead of pure collision avoidance. Here, we reevaluate the case where the robot overtakes Alice, who does not see the robot behind her. This time, we increase the sizes and from 0.5 m to 1.0 m. Correspondingly, Eq. 2 will result in a conflict whenever m. Assuming circular agents with a diameter of 0.5 m, this amounts to 1 m of free space between the agents, which can roughly be considered the boundary of personal space [13]. Figure 9 shows the resulting Responsibility and Engagement, evaluated with respect to Collision Avoidance and Personal Space Compliance. By passing close to Alice, the robot avoids a collision, but violates her personal space. Accordingly, using the increased size, the robot’s Responsibility is now lower, with . Furthermore, there is a residual conflict at the time of closest encounter, which is captured by . Since the robot did not actively turn towards Alice after evading the collision, the corresponding Engagement is still zero ().
@@ V-A Research Questions, Interpretability, and Grounding
With respect to our first research question, the dyadic interaction scenarios of Experiments 1 and 4 show that the proposed metrics reliably quantify which agent contribute to escalating or diminishing a conflict, in a variety of different geometric constellations. Furthermore, Experiment 5 shows that they also capture residual conflicts caused by a collision avoidance algorithm entering the personal space of another agent.
@@
The group splitting scenario of addresses the second research question. Here, the different trajectory options for the robot share the same evaluation context (e.g., a similar scenario start and progression). In this case, the values of and enable a direct comparison of different trajectory alternatives (going left, right, or through the group). The ordinal ranking of the metric matches a subjective ranking of the behavior quality.

### v2

@@ I Introduction
One goal of Social Robot Navigation (SRN) research is to enable robots to seamlessly navigate complex, unstructured, and potentially crowded environments. Interactions between robots and humans are at the core of SRN. Studies show that humans interacting with publicly deployed robots often compensate for the robots’ lack of navigation capabilities, thereby taking a major share of preventing collisions and conflicting situation [1, 11]. While this might be acceptable for the occasional robot, it could turn out to be a nuisance once more robots are deployed in public. Therefore, research on socially acceptable planning algorithms as well as suitable evaluation metrics has become increasingly important [15, 5, 4]. To assess the quality of a navigation algorithm, a meaningful quantitative evaluation of the output trajectories is paramount. Such metrics can also be part of target functions for learning-based approaches. In this work, we focus on metrics which evaluate various aspects of social compliance of motion behavior, i.e., trajectories. In [16], we pointed out that current metrics fail to address the causal component in interactions between two agents. The degree to which an agent contributed to resolving the conflict is vital for evaluating the quality of said agent’s behavior. To this end, we proposed a ”Responsibility” metric aiming to quantify the share that agents contributed to de-escalating a potential conflict.
@@ II Related work
In recent work [16], we showed that frequently used existing metrics do not capture important aspects of two moving agents interacting with each other. Specifically, we analyzed kinematics-based metrics (e.g. an agent’s maximum acceleration or mean jerk), metrics based on the distance between two agents (e.g., Clearing Distance or Space Compliance Rate) and prediction-based metrics like Minimum Time-To-Collision or Projected Path Duration [9, 15] on a simple scenario: two agents of varying capabilities (oblivious or foresighted) are facing each other on a head-on collision course. While all of the established metrics show meaningful information, they fail to fully highlight a key information: which of the agents (if any) were causally responsible for avoiding the conflict. [16] propose the Responsibility metric, which is based on predicted minimal distances. In the presented scenarios, the Responsibility metric shows which agent’s behavior contributed to the relaxation of the conflict. However, the set of tested scenarios is limited, and all scenarios start with both agents being at full conflict already. Furthermore, no scenarios with more than two agents are addressed.
Other work has also addressed the topic of responsibility in the context of navigation explicitly. [6] captures causal responsibility of moving agents in their proposed FeAR metric in a discrete, grid-based representation, also assuming that attribution of responsibility should be related to actions (or their lack of). Other approaches from the automotive domain also capture or estimate in-situ responsibility, often relying on strong geometrical priors such as lanes or regions [12, 13]. Here, we focus on the continuous case, in mostly unstructured environments. We strive for a simple metric for ex-post trajectory evaluation. Note that the concept of responsibility itself is widely used in many different contexts (philosophical, legal, technical, …), but, at the same time, inherently ambiguous [14]. This paper focuses on outcome or causal responsibility responsibility in [14]’s taxonomy.
@@ III Responsibility and Engagement Metrics
This section extends the Responsibility metric presented in [16]. First, we revisit the calculation of the underlying conflict potential in Section III-A. We then propose to derive the conflict measure by considering the previously unaddressed conflict buildup phase using a normalization step in Section III-B. Section III-C explains how the agents’ actions are taken into consideration for the escalation or de-escalation of the conflict, before Section III-D shows the updated Responsibility metric and introduces the Engagement metric. Program code for calculating the Responsibility and Engagement metrics from trajectories can be found in a GitHub repository111https://github.com/HRI-EU/hri-metrics.
@@ III-A Conflict Potential
We use the definition of [16] to calculate the Conflict Potential between two agents (referred to as agent 1 and agent 2) in an interaction. Here, the Distance of Closest Encounter (DCE) [2] of two trajectories plays a central role. To calculate the Conflict Potential, a predictive DCE (pDCE) is used, which captures the predicted proximity of two agents continuing along their current trajectories with constant velocity. Figure 1 shows how the pDCE captures the minimum distance between two agents when their movements are predicted based on a constant velocity assumption.
@@
where and are the radii of agents 1 and 2, assuming circular agents. In a nutshell, the conflict potential equals one if the agents are on a direct collision course. It decreases as their predicted center positions at the point of closest encounter move further apart, becoming zero when they are predicted to just miss each other narrowly.
@@ III-B Conflict
Figure 2 illustrates the relationship between conflict potential, the time normalization and the resulting conflict measure. See Section V for a discussion on alternatives to using the TCE and the choice of the window size .
@@ III-C Agent Behavior and Conflict Contribution
One interpretation is that is the residual conflict which was not fully resolved by the agents. As the distance between agents increases again, any remaining conflict is therefore diminished by time.
@@ IV-B Experiment 1: Dyadic Interactions
This experiment provides two insights: First, since the Responsibility metric is based on the pDCE, it covers diverse interactions, regardless of the orientation of both agents toward each other. Also, it captures both longitudinal and lateral behavior changes. Second, while Responsibility is a valuable metric to investigate behavior in interactions, it does not entail any direct normative interpretation. In order to evaluate if a behavior was acceptable, the context is vital. Consider Scenarios 2 and 3, where only one agent is responsive, for the Oncoming and Overtaking case. Normatively, we would expect an agent to be alert, taking others into account when planning its trajectory. Hence, a Responsibility of zero typically corresponds to suboptimal (or at least unfriendly) behavior. However, in Scenario 3 of the Overtaking geometry, Alice’s oblivious behavior is fully acceptable, since she can’t be expected to see the robot approaching from behind. Vice versa, in Scenario 2, the robot is threatening to run over Alice from behind - which is completely unacceptable, since Alice has to jump away to save herself. Similarly, in the Overtaking case, Alice’s Engagement would rise if she slowed down while the robot is approaching from behind (since she actively changes her behavior).
@@ IV-E Experiment 4: Playing Catch
Figure 8 shows one interaction of the game. The robot starts on the bottom right corner going diagonally to the upper left corner, with Alice starting in the bottom left corner, tying to intercept its way. The robot reacts by evading to the left. Alice readjusts the course but overshoots, starting to go in a narrow right curve, just avoiding a collision. the robot keeps evading. Since this point in time marks the interaction’s time of closest encounter (TCE, see Eq. 3), the rest of the interaction (i.e., Alice and the robot completing the circle) is not captured.
@@ IV-F Experiment 5 - Personal Space Compliance
In Eq. 2, we can choose the agent sizes to determine which predicted distance leads to a potential conflict. Accordingly, it is possible to evaluate and with respect to proxemics [7] instead of pure collision avoidance. Here, we reevaluate the case where the robot overtakes Alice, who does not see the robot behind her. This time, we increase the sizes and from 0.5 m to 1.0 m. Correspondingly, Eq. 2 will result in a conflict whenever m. Assuming circular agents with a radius of 0.5 m, this amounts to 1 m of free space between the agents, which can roughly be considered the boundary of personal space [7]. Figure 9 shows the resulting Responsibility and Engagement, evaluated with respect to Collision Avoidance and Personal Space Compliance. By passing close to Alice, the robot avoids a collision, but violates her personal space. Accordingly, using the increased size, the robot’s Responsibility is now lower, with . Furthermore, there is a residual conflict at the time of closest encounter, which is captured by . Since the robot did not actively turn towards Alice after evading the collision, the corresponding Engagement is still zero ().
@@ V-A Research Questions, Interpretability, and Grounding
With respect to our first research question, the dyadic interaction scenarios of Experiments 1 and 4 show that the proposed metrics reliably quantify which agent contribute to escalating or diminishing a conflict, in a variety of different geometric constellations, thereby complementing existing metrics [16]. Furthermore, Experiment 5 shows that they also capture residual conflicts caused by a collision avoidance algorithm entering the personal space of another agent.
@@
The group splitting scenario of Experiment 2 addresses the second research question. Here, the different trajectory options for the robot share the same evaluation context (e.g., a similar scenario start and progression). In this case, the values of and enable a direct comparison of different trajectory alternatives (going left, right, or through the group). The ordinal ranking of the metric matches a subjective ranking of the behavior quality.